The lengths of mature trout in a local lake are approximately normally distributed with a mean of 12.1 inches, and a standard deviation of 1.6 inches.
μ=12.1

σ=1.6
Fill in the indicated boxes.
Find the z-score corresponding to a fish that is 12.9 inches long. Round your answer to the nearest hundredth as needed.

How many inches long is a fish that has a z-score of 1? Round your answer to the nearest tenth as needed.

Answer by Boreal(15235)

z=(x-mean)/sd
=(12.9-12.1)/1.6
=+0.5
--
1=(x-12.9)/16.
1.6=x-12.9
x=14.5 inches